# Unclonable Encryption with Continuous Variables

### v1

@@ 1 Introduction
The term Unclonable Encryption (UE) first appeared in 2003 in a paper by Gottesman [Gottesman2003]. Alice encrypts a classical message into a quantum state. A security definition was introduced that essentially states “If Bob decides that his decryption is valid, then Eve, given the key, has only negligible information about the plaintext.” The security definition was formulated in terms of the trace distance between encryptions of different messages. In the same framework, Leermakers and Škorić devised an UE scheme with key recycling [Leermakers2021]. Broadbent and Lord [broadbent_et_al2020] introduced a modified security definition, based on a cloning game, and constructed UE in the random oracle model. Several further UE schemes were introduced in [Ananth2021, Leermakers2022, KunduTan2025], and results on the feasibility and limitations of UE were given in [Ananth2022, Majenz2021].
@@
Until now, UE has been exclusively studied in discrete-variable (DV) quantum systems. However, in the field of Quantum Key Distribution (QKD), continuous-variable (CV) quantum systems have emerged as an attractive alternative to DV [PhysRevA.61.010303, PhysRevA.61.022309, PhysRevA.62.062308, PhysRevA.63.052311, PhysRevLett.88.057902, article_Grossmman_03, 10.5555/2011564.2011570, PhysRevLett.93.170504, garcia2007quantum, leverrier:tel-00451021]. CV does not need expensive single-photon detectors, and has the advantage that low-loss telecom wavelengths (1310nm, 1550nm) can be used, making it possible to capitalize on several decades of experience in coherent optical communication technology. Beyond QKD, the practical advantages hold more generally for other quantum information processing applications, and this has fueled substantial interest in translating DV-based cryptographic ideas into the CV domain.
@@
In this paper, we propose the first Unclonable Encryption scheme that works with Continuous-Variable states. We provide a security proof in the UE framework of [broadbent_et_al2020]. It turns out that bringing UE from discrete to continuous variables has a number of nontrivial aspects. On the construction side, the parameters of the scheme need to be tuned such that there both decryptability and unclonability are satisfied. On the proof-technical side, we introduce a number of ‘game hops’ in order to connect the cloning game, which features in the UE security definition, to the CV monogamy-of-entanglement game, for which an upper bound on the winning probability has recently been proven [Culf2022]. Furthermore it is necessary to slightly modify the definition of an encryption, in order to allow for a small probability of decryption failure, caused by the probability tails of the CV measurements.
@@
The outline of the paper is as follows. In Section 2 we briefly review CV formalism and important definitions and results from the literature. In Section 3 we present our protocol and verify that it satisfies the definition of a quantum encryption. Section 4 contains the unclonability proof. The main result is stated in Theorem 4.3. In Section 5 we summarize and discuss future work.
@@ 2.1 Notation
We use standard bra-ket notation for quantum states. Hilbert spaces are written as ${\cal H}$ with a subscript. E.g. we write the Hilbert space of a single CV mode as ${\cal H}_{1}$. The notation ${\cal D}({\cal H})$ stands for the set of density operators on ${\cal H}$. The Hamming weight of a string $s$ is written as $|s|$.
@@ 2.2 Continuous Variables; Gaussian states
A ‘mode’ of the electromagnetic vector potential represents a plane wave solution of the vacuum Maxwell equations at a certain frequency, wave vector and polarisation. Associated with each mode there is a creation operator $\hat{a}^{\dagger}$ and annihilation operator $\hat{a}$. The linear combinations $\hat{q}=\frac{\hat{a}+\hat{a}^{\dagger}}{\sqrt{2}}$ and $\hat{p}=\frac{\hat{a}-\hat{a}^{\dagger}}{i\sqrt{2}}$ are easy-to-observe quantities called quadratures, and they behave as the position and momentum operator of a harmonic oscillator.
[…]
@@ 2.3 Definitions and useful lemmas
$\mathrm{Key}_{\lambda}$ $:\mathcal{D}(\mathcal{C})\to\mathcal{D}(\mathcal{H}_{{\cal K}(\lambda)})$ (1)
$\mathrm{Enc}_{\lambda}$ $:\mathcal{D}(\mathcal{H}_{{\cal K}(\lambda)}\otimes\mathcal{H}_{{\cal M}(\lambda)})\to\mathcal{D}(\mathcal{H}_{T(\lambda)})$ (2)
$\mathrm{Dec}_{\lambda}$ $:\mathcal{D}(\mathcal{H}_{{\cal K}(\lambda)}\otimes\mathcal{H}_{T(\lambda)})\to\mathcal{D}(\mathcal{H}_{{\cal M}(\lambda)})$ (3)
@@
Unclonability of a QECM scheme is defined via a cloning game. A challenger prepares a cipherstate and gives it to Alice. Alice splits the cipherstate into two pieces; one piece goes to Bob, one to Charly. Then Bob and Charly receive the decryption key and must both produce the correct plaintext, without being allowed to communicate. A QECM scheme is considered to be secure against cloning if the three players ABC, acting together, have an exponentially small probability of winning the game.
We follow the definitions of [broadbent_et_al2020] for the cloning game and the security against cloning.
@@
Our security proof will make use of recent results on entanglement monogamy games for CV systems [Culf2022]. Such a game is played between Alice on one side and Bob and Charlie on the other. Bob and Charlie prepare a tripartite state $\rho_{ABC}$ of their choice. Then Alice does an unpredictable measurement, and Bob and Charlie have to show that they are both sufficiently entangled with Alice to guess Alice’s outcome to some degree of accuracy. Below we present the coset monogamy game on $\mathbb{R}^{N}$ as introduced in Section 4 of [Culf2022]. In the original version, the measurement bases are continuous variable coset states of $\mathbb{R}^{N}$, given by $|P_{q,p}\rangle=\bigotimes_{i=1}^{N}\begin{cases}|q=q_{i}\rangle,&i\notin I,\\ |p=p_{i}\rangle,&i\in I.\end{cases}$ Here, $I$ is a subset interval of $N$. We note that this is equivalent to performing a homodyne measurement in the $q$ direction, if $i\in I$, otherwise in the $p$ direction. We refer to this game as the ‘partial’ CV monogamy of entanglement game, since it requires Bob and Charlie to produce guesses for only one quadrature direction instead of two. In Section 4.4 we introduce a ‘full’ version of the game.
@@ 3 Protocol Description
For proof-technical reasons, the squeezing directions are not chosen uniformly at random. We impose the constraint that exactly half the modes are squeezed in the $q$-direction, and one half in the $p$-direction. Hence the corresponding key space is not $\{0,1\}^{N}$ but rather the set =${\cal L}=\{1,\ldots,\log_{2}{N\choose N/2}\}$ of labels which uniquely enumerate the strings with Hamming weight $N/2$.
@@
Our construction is a QECM scheme with decryption failure parameter $\varepsilon_{\rm DF}=\exp\Big{[}-ND_{\rm KL}(\frac{t+1}{N}||\frac{1}{2}{\rm Erfc}(\alpha\sqrt{\cosh r}))\Big{]}$, where $D_{\rm KL}$ stands for the binary Kullback-Leibler divergence, $D_{\rm KL}(a||b)=a\ln\frac{a}{b}+(1-a)\ln\frac{1-a}{1-b}$, and Erfc is the complementary error function.
@@
It is trivial to see that the triplet (Key, Enc, Dec) fits the format of the CPTP maps in Def. 2.1. All that is left to show is that the protocol satisfies the correctness condition. A bit error in the codeword bit $c_{i}$ occurs when $y_{i}-k_{i}$ has the wrong sign. Consider, without loss of generality, $c_{i}=0$. Then the bit error probability is $\beta:={\rm Pr}[Y_{i}-k_{i}<0]$. When $k_{i}$ is known, the random variable $Y_{i}-k_{i}$ is gaussian-distributed with mean $\alpha$ and variance $1/(2\cosh r)$. Hence $\beta=\frac{1}{2}-\frac{1}{2}{\rm Erf}(\alpha\sqrt{\cosh r})$.
A decoding error can occur only when there are more than $t$ bitflips; this occurs with probability $P=\sum_{j=t+1}^{N}{N\choose j}\beta^{j}(1-\beta)^{N-j}$ $=\sum_{\ell=0}^{N-t-1}{N\choose\ell}\beta^{N-\ell}(1-\beta)^{\ell}$. Using a Chernoff bound on the binomial tail, we obtain
$P\leq\exp\Big{[}-ND_{\rm KL}(\frac{t+1}{N}||\beta)\Big{]}.$ (11)
As an illustration, consider a ciphertext of size $N=1000$. We set the protocol parameters $\alpha=0.05$ and $r=8$. This leads to a bit error probability of $\beta=0.00317$. We choose $t=16(>N\beta=3.16)$. This leads to an exponentially small failure probability $\varepsilon_{\rm DF}=3.63\times 10^{-7}$.
@@ 4.1 Game hopping
We prove that our scheme satisfies the security definition of Def. 2.3. We do this as follows. (i) We rewrite the state preparation from the original prepare-and-send form to Entanglement Based (EB) form. (ii) We show, in a number of ‘hops’, that the cloning game that the security definition is based on is equivalent to a CV entanglement monogamy game. (iii) Finally we use Lemma 2.5 which upper bounds the winning probability in the CV entanglement monogamy game.
@@ 4.3 Unclonable Encryption game
Prepare-and-send: $Ch$ runs Key (algorithm 1) and obtains $(s,\phi,k)$. He runs Enc (algorithm 2) on $m$. He sends the resulting cipherstate to Alice.
Entanglement-based: $Ch$ runs part of algorithm 1 and obtains $(s,\phi)$. He prepares an entangled state $\rho_{ChA}=\bigotimes_{i=1}^{N}\rho_{ChA}^{(i)}$, as explained in Section 4.2. He sends the ‘A’ part to Alice.
@@
Prepare-and-send: $Ch$ announces the key $(s,\phi,k)$. $\mathcal{B}$ and $\mathcal{C}$ respond with $m_{B}$ and $m_{C}$ respectively.
Entanglement-based: $Ch$ does homodyne measurements on his modes and computes $k$ from the outcomes. $Ch$ announces the key $(s,\phi,k)$. $\mathcal{B}$ and $\mathcal{C}$ respond with $m_{B}$ and $m_{C}$ respectively.
@@
$Ch$ prepares the entangled state $\rho_{ChA}$. He sends one half of the state to Alice.
@@
$Ch$ measures $u$ and announces $\phi$. ${\cal B}$ and ${\cal C}$ respond with $u_{B}$ and $u_{C}$ respectively. (Their guess for $u$.)
@@
For $i\in\{1,\ldots,N\}$ $Ch$ calculates $c_{Bi}=\frac{1}{2}-\frac{1}{2}{\rm sign}(u_{Bi}-\frac{k_{i}}{\tanh r})$ and $c_{Ci}=\frac{1}{2}-\frac{1}{2}{\rm sign}(u_{Ci}-\frac{k_{i}}{\tanh r})$. Then $m_{B}={\tt Dec}_{\rm base}(s,{\tt Decode}(c_{B}))$ and $m_{C}={\tt Dec}_{\rm base}(s,{\tt Decode}(c_{C}))$. The players ${\cal A}{\cal B}{\cal C}$ win if $m_{B}=m_{C}=m$.

### v2

# Practical Unclonable Encryption with Continuous Variables

@@ Abstract
Furthermore, we demonstrate that our scheme can be readily implemented with current technology. By incorporating realistic imperfections such as channel noise and detector inefficiencies, we show that the protocol remains robust under these conditions.
@@ 1 Introduction
The term Unclonable Encryption (UE) first appeared in 2003 in a paper by Gottesman [Gottesman2003]. Alice encrypts a classical message into a quantum state. A security definition was introduced that essentially states “If Bob decides that his decryption is valid, then Eve, given the key, has only negligible information about the plaintext.” The security definition was formulated in terms of the trace distance between encryptions of different messages, and essentially encompasses tamper detection. In the same framework, Leermakers and Škorić devised an UE scheme with key recycling [Leermakers2021]. Broadbent and Lord [broadbent_et_al2020] introduced the stronger notions of unclonability and unclonable indistinguishability, in which the splitting of the ciphertext is considered, and both receiving partners are malicious. They constrructed UE in the random oracle model. Several further UE schemes were introduced in [Ananth2021, Leermakers2022, KunduTan2025], and results on the feasibility and limitations of UE were given in [Ananth2022, Majenz2021].
@@
Until now, UE has been exclusively studied in discrete-variable (DV) quantum systems. However, in the field of Quantum Key Distribution (QKD), continuous-variable (CV) quantum systems have emerged as an attractive alternative to DV [PhysRevA.61.010303, PhysRevA.61.022309, PhysRevA.62.062308, PhysRevA.63.052311, PhysRevLett.88.057902, article_Grossmman_03, 10.5555/2011564.2011570, PhysRevLett.93.170504, garcia2007quantum, leverrier:tel-00451021]. A key advantage of CV-QKD over its DV equivalent is its practicality, see, e.g., Ref. [RevModPhys.84.621]. In broad terms, CV platforms are simpler to operate and exploit decades of progress in coherent optical-communication technology. DV protocols, by contrast, need single-photon detectors and sometimes single-photon sources as well. These components remain expensive and technically demanding, especially when photon-number resolution is required. CV schemes instead use homodyne or heterodyne detection, which is both cheaper and easier to deploy. Furthermore, today’s fibre-optic infrastructure is optimised for low-loss transmission at telecom wavelengths (1310 nm and 1550 nm); ideal single-photon emitters at these wavelengths have not yet been realised, and frequency up-conversion introduces additional loss and error. Beyond QKD, the practical advantages hold more generally for other quantum information processing applications, and this has fueled substantial interest in translating DV-based cryptographic ideas into the CV domain.
In this paper, we propose the first Unclonable Encryption scheme that works with Continuous-Variable states. We prove that our scheme is unclonable secure, in the UE framework of [broadbent_et_al2020]. We note that this is strictly weaker than the more general notion of unclonable indistinguishability. Nevertheless, for the practical objective of a one-time encryption scheme, unclonable security already ensures that no adversary can duplicate or replay ciphertexts without detection, which suffices for our application. Furthermore, unclonable indistinguishability allows for the adversary to be entangled with the plaintext, which gives too much power to the adversary, since the plaintext is classical.
@@
It turns out that bringing UE from discrete to continuous variables has a number of nontrivial aspects. On the construction side, the parameters of the scheme need to be tuned such that both decryptability and unclonability are satisfied. On the proof-technical side, we use a series of hybrids in order to connect the cloning game, which features in the UE security definition, to the CV monogamy-of-entanglement game, for which an upper bound on the winning probability has recently been proven [Culf2022]. Furthermore it is necessary to allow imperfect completeness in the CV case, as there is a non-zero probability of decryption failure caused by the probability tails of the CV measurements.
@@
In line with this theme of practicality, we also show that our scheme is correct for experimentally feasible values of squeezing. Finally we analyse our protocol under noisy conditions and show robustness for realistic channel parameters.
The outline of the paper is as follows. In Section 2 we briefly review CV formalism and important definitions and results from the literature. In Section 3 we present our protocol and verify that it satisfies the definition of a quantum encryption. Section 4 contains the unclonability proof. The main result is stated in Theorem 4.4. In section 5 we discuss squeezing and robustness under channel/detector imperfections. In Section 6 we summarize and discuss future work.
@@ 2.1 Notation
[…]
@@ 2.2 Continuous Variables; Gaussian states
A ‘mode’ of the electromagnetic vector potential represents a plane wave solution of the vacuum Maxwell equations at a certain frequency, wave vector and polarisation. Associated with each mode there is a creation operator $\hat{a}^{\dagger}$ and annihilation operator $\hat{a}$. The linear combinations $\hat{q}=\frac{\hat{a}+\hat{a}^{\dagger}}{\sqrt{2}}$ and $\hat{p}=\frac{\hat{a}-\hat{a}^{\dagger}}{i\sqrt{2}}$ are easy-to-observe quantities called quadratures, and they behave as the position and momentum operator of a harmonic oscillator. Measuring a single quadrature is called a homodyne measurement.
The Gaussian states are a special class of CV states; their Wigner function (quasi density function on the phase space) [RevModPhys.84.621] is a Gaussian function of the quadrature variables. An $N$-mode Gaussian state is fully characterized by a displacement vector $d\in\mathbb{R}^{2N}$ and $2N\times 2N$ covariance matrix $\Gamma$. The corresponding Wigner function is $\frac{1}{\pi^{N}\sqrt{\det\Gamma}}\exp-(x-d)^{T}\Gamma^{-1}(x-d)$, where $x$ stands for the vector $(q_{1},p_{1},\ldots,q_{N},p_{N})^{T}$. The class of Gaussian states contains important states like the vacuum, thermal states, coherent states, squeezed coherent states and EPR states (two-mode squeezed vacuum). A coherent state that is squeezed in the $q$-direction has covariance matrix $e^{-\zeta}\;0\;\choose 0\;\;\;\;e^{\zeta}$}, where $\zeta\geq 0$ is the squeezing parameter. For the $p$-direction it is $e^{\zeta}\;\;0\;\choose 0\;\;e^{-\zeta}$}. All intermediate directions are possible, but will not be used in this paper.
An EPR state is obtained by mixing a $q$-squeezed vacuum with a $p$-squeezed vacuum using a 50/50 beam splitter. The resulting two-mode squeezed (TMS) state has zero displacement, and its covariance matrix is $\mathbb{I}\cosh\zeta\;\;\;\;\sigma_{z}\sinh\zeta\choose\sigma_{z}\sinh\zeta\;\;\;\mathbb{I}\cosh\zeta$}, where $\mathbb{I}$ is the $2\times 2$ identity matrix and $\sigma_{z}={1\;\;\;0\choose 0\;-1}$. In the entanglement based version of our scheme we will make use of a state that resembles a displaced EPR state, with for instance displacement $d=(\alpha,0,\alpha,0)^{T}$. When the $q$-quadrature is measured on one side, the probability density for measurement outcome $x_{A}$ is a normal distribution with mean $\alpha$ and variance $\frac{1}{2}\cosh\zeta$. The measurement projects the state on the other side to a squeezed coherent state with displacement $(\alpha+[x_{A}-\alpha]\tanh\zeta,0)^{T}$ and covariance matrix $\frac{1}{\cosh\zeta}\;\;\;0\;\choose 0\;\;\;\;\cosh\zeta$}. If the $x_{A}$ is not known, the displacement $\alpha+[x_{A}-\alpha]\tanh\zeta$ is a stochastic variable following a Gaussian distribution with mean $\alpha$ and variance $\frac{1}{2}\cosh\zeta\tanh^{2}\zeta$.
@@ 2.3 Definitions and useful lemmas
$\text{Key}_{\lambda}$ $:\mathcal{D}(\mathcal{C})\to\mathcal{D}(\mathcal{H}_{{\cal K}(\lambda)})$ (1)
$\text{Enc}_{\lambda}$ $:\mathcal{D}(\mathcal{H}_{{\cal K}(\lambda)}\otimes\mathcal{H}_{{\cal M}(\lambda)})\to\mathcal{D}(\mathcal{H}_{T,\lambda})$ (2)
$\text{Dec}_{\lambda}$ $:\mathcal{D}(\mathcal{H}_{{\cal K}(\lambda)}\otimes\mathcal{H}_{T,\lambda})\to\mathcal{D}(\mathcal{H}_{{\cal M}(\lambda)})$ (3)
@@
Unclonability of a QECM scheme is defined via a cloning game. A challenger prepares a cipherstate and gives it to Alice. Alice splits the cipherstate into two pieces; one piece goes to Bob, one to Charlie. Then Bob and Charlie receive the decryption key and must both produce the correct plaintext, without being allowed to communicate. A QECM scheme is considered to be secure against cloning if the three players ABC, acting together, have an exponentially small probability of winning the game. We follow the definitions of [broadbent_et_al2020] for the cloning game and the security against cloning.
@@
Our security proof will make use of recent results on entanglement monogamy games for CV systems [Culf2022]. Such a game is played between Alice on one side and Bob and Charlie on the other. Bob and Charlie prepare a tripartite state $\rho_{ABC}$ of their choice. Then Alice does an unpredictable measurement, and Bob and Charlie have to show that they are both sufficiently entangled with Alice to guess Alice’s outcome to some degree of accuracy. Below we present the coset monogamy game on $\mathbb{R}^{N}$ as introduced in Section 4 of [Culf2022]. In the original version, the measurement bases are continuous variable coset states of $\mathbb{R}^{N}$, given by $|P_{q,p}\rangle=\bigotimes_{i=1}^{N}\begin{cases}|q=q_{i}\rangle,&i\notin I,\\ |p=p_{i}\rangle,&i\in I.\end{cases}$ Here, $I$ is a subset interval of $N$. We note that this is equivalent to performing a homodyne measurement in the $q$ direction if $i\in I$, otherwise in the $p$ direction. We refer to this game as the ‘partial’ CV monogamy of entanglement game, since it requires Bob and Charlie to produce guesses for only one quadrature direction instead of two. In Section 4.4 we introduce a ‘full’ version of the game.
@@ 3 Protocol Description
For proof-technical reasons, the squeezing directions are not chosen uniformly at random. We impose the constraint that exactly half the modes are squeezed in the $q$-direction, and one half in the $p$-direction. Hence the corresponding key space is not $\{0,1\}^{N}$ but rather the set ${\cal L}=\{1,\ldots,\log_{2}{N\choose N/2}\}$ of labels which uniquely enumerate the strings with Hamming weight $N/2$.
@@
Our construction is a QECM scheme with decryption failure parameter
$\varepsilon_{\rm DF}=\exp\Big[-ND_{\rm KL}(\frac{t+1}{N}||\frac{1}{2}{\rm Erfc}(\alpha\sqrt{\cosh r}))\Big],$ (11)
where $D_{\rm KL}$ stands for the binary Kullback-Leibler divergence, $D_{\rm KL}(a||b)=a\ln\frac{a}{b}+(1-a)\ln\frac{1-a}{1-b}$, and Erfc is the complementary error function.
@@
It is trivial to see that the triplet (Key, Enc, Dec) fits the format of the CPTP maps in Def. 2.1. All that is left to show is that the protocol satisfies the correctness condition. A bit error in the codeword bit $c_{i}$ occurs when $y_{i}-k_{i}$ has the wrong sign. Consider, without loss of generality, $c_{i}=0$. Then the bit error probability is $\beta:={\rm Pr}[Y_{i}-k_{i}<0]$. When $k_{i}$ is known, the random variable $Y_{i}-k_{i}$ is gaussian-distributed with mean $\alpha$ and variance $1/(2\cosh r)$. Hence
$\beta=\frac{1}{2}-\frac{1}{2}{\rm Erf}(\alpha\sqrt{\cosh r}).$ (12)
A decoding error can occur only when there are more than $t$ bitflips; this occurs with probability $P=\sum_{j=t+1}^{N}{N\choose j}\beta^{j}(1-\beta)^{N-j}$ $\sum_{\ell=0}^{N-t-1}{N\choose\ell}\beta^{N-\ell}(1-\beta)^{\ell}$. Using a Chernoff bound on the binomial tail, we obtain $P\leq\exp\Big[-ND_{\rm KL}(\frac{t+1}{N}||\beta)\Big]$. ∎
Note that asymptotically $N$ gets close to $\frac{n}{1-h(\beta)}$, where $h$ is the binary entropy function.
@@ 4.1 Game hopping
We prove that our scheme satisfies the security definition of Def. 2.3. We do this as follows. (i) We rewrite the state preparation from the original prepare-and-send form to Entanglement Based (EB) form. (ii) We show, in a series of ‘hops’ (hybrids), that the cloning game that the security definition is based on is equivalent to a CV entanglement monogamy game. (iii) Finally we use Lemma 2.5 which upper bounds the winning probability in the CV entanglement monogamy game.
@@ 4.3 Unclonable Encryption game
Prepare-and-send: ChCh runs Key (algorithm 1) and obtains $(s,\phi,k)$. He runs Enc (algorithm 2) on $m$. He sends the resulting cipherstate to Alice.
Entanglement-based: ChCh runs part of algorithm 1 and obtains $(s,\phi)$. He prepares an entangled state $\rho_{ChA}=\bigotimes_{i=1}^{N}\rho_{ChA}^{(i)}$, as explained in Section 4.2. He sends the ‘A’ part to Alice.
@@
Prepare-and-send: ChCh announces the key $(s,\phi,k)$. $\mathcal{B}$ and $\mathcal{C}$ respond with $m_{B}$ and $m_{C}$ respectively.
Entanglement-based: ChCh does homodyne measurements on his modes and computes $k$ from the outcomes. ChCh announces the key $(s,\phi,k)$. $\mathcal{B}$ and $\mathcal{C}$ respond with $m_{B}$ and $m_{C}$ respectively.
@@
ChCh prepares the entangled state $\rho_{ChA}$. He sends one half of the state to Alice.
@@
[…]
@@
For $i\in\{1,\ldots,N\}$ ChCh calculates $c_{Bi}=\frac{1}{2}-\frac{1}{2}{\rm sign}(u_{Bi}-\frac{k_{i}}{\tanh r})$ and $c_{Ci}=\frac{1}{2}-\frac{1}{2}{\rm sign}(u_{Ci}-\frac{k_{i}}{\tanh r})$. Then $m_{B}={\tt Dec}_{\rm base}(s,{\tt Decode}(c_{B}))$ and $m_{C}={\tt Dec}_{\rm base}(s,{\tt Decode}(c_{C}))$. The players ${\cal A}{\cal B}{\cal C}$ win if $m_{B}=m_{C}=m$.